OpenStudy (anonymous):

Solve using the substitution method 3x + 2y = 3, 9x − 8y = −2

OpenStudy (anonymous):

well what do u know about complementary angles?

OpenStudy (anonymous):

So, the two angles together are 90 degrees. Can you turn the info you have into an equation?

OpenStudy (anonymous):

Well the Question that is being asked is that what are the two complementary angles.... and they give us that one part equals4x-2 and the other one equals x

OpenStudy (anonymous):

and beacuse when you add complementary angles it equals to 90

OpenStudy (anonymous):

Your equation should look like

OpenStudy (anonymous):

Liza2, yes I just don't know how to brake it down.... I have so far this: <1=4x-2 <2=x <1+<2=90 deg and from there am stuck

OpenStudy (anonymous):

\[4x-2 + x =90\]

OpenStudy (anonymous):

I dont know what to do next

OpenStudy (anonymous):

Can you isolate "x" using the equation @CouldntFindAGoodName just gave you? 4x - 2 + x = 90 Because once you've isolated x, you can figure out the measurements of both angles

OpenStudy (anonymous):

Both angles equal 90 so u just add the angles and get 4x-2 +x =90

OpenStudy (anonymous):

Then you isolate the x

OpenStudy (anonymous):

But beforew you isolate the x u add like terms

OpenStudy (anonymous):

SO it would be like 4x+x -2=90 which equals ---> 4x-2 =90

OpenStudy (anonymous):

am confused, so how would I add, is it like this: x+(4x-2)=90 5x-2=90

OpenStudy (anonymous):

* 5x-2=90

OpenStudy (anonymous):

Then you isolate the x

OpenStudy (anonymous):

awesome I have that so far CouldntFindAGoodName

OpenStudy (anonymous):

First you add 2 to both sides of the equation

OpenStudy (anonymous):

how would I isolate the x thou?

OpenStudy (anonymous):

First add 2 to both side of the equation

OpenStudy (anonymous):

like (5x-2) +2 =90 +2 ----> to get 5x=92

OpenStudy (anonymous):

Then from there to isolate the x you divide the 5 on both side like (5x)/5=92/5

OpenStudy (anonymous):

then u are left with x=18.4

OpenStudy (anonymous):

and for your x value would equal 18.4

OpenStudy (anonymous):

then you just plug in your x value for your two complementary angles

OpenStudy (anonymous):

1< 4x-2 ---> 4(18.4)-2 2< x----> 18.4

OpenStudy (anonymous):

xD no problem XD